Jurvin Enterprises is a manufacturing company that had no beginning inventories. A subset of the transactions that it recorded during a recent month is shown below.

a. $94,000 in raw materials were purchased for cash.

b. $89,000 in raw materials were used in production. Of this amount, $78,000 was for direct materials and the remainder was for indirect materials..

c. Total labor wages of $132,000 were incurred and paid. Of this amount, $112,000 was for direct labor and the remainder was for indirect labor.

d. Additional manufacturing overhead costs of $143,000 were incurred and paid.

e. Manufacturing overhead of $152,000 was applied to production using the company’s predetermined overhead rate.

f. All of the jobs in process at the end of the month were completed.

g. All of the completed jobs were shipped to customers.

h. Any underapplied or overapplied overhead for the period was closed to Cost of Goods Sold.

Required:

1. Post the above transactions to T-accounts.

2. Determine the adjusted cost of goods sold for the period.

  • Required 1
  • Required 2

Post the above transactions to T-accounts.

Cash Raw Materials
Beg. Bal. Beg. Bal.
a.
b.
c.
End. Bal. End. Bal.
Work in Process Finished Goods
Beg. Bal. Beg. Bal.
End. Bal. End. Bal.
Manufacturing Overhead Cost of Goods Sold
Beg. Bal. Beg. Bal.
End. Bal. End. Bal.
  • Required 1
Answer #1

Required 1

Cash    Raw Materials
Beg. Bal. a. 94000 Beg. Bal.
c. 132000 a. 94000 b. 89000
d. 143000
End. Bal. $369000 End. Bal. $5000
Work in Process Finished Goods
Beg. Bal. Beg. Bal.
b. 78000 f. 342000 f. 342000 g. 342000
c. 112000
e. 152000
End. Bal. $0 End. Bal. $0
Manufacturing Overhead Cost of Goods Sold
Beg. Bal. Beg. Bal.
b. 11000 e. 152000 g 342000
c. 20000 h. 22000 h 22000
d. 143000
End. Bal. $0 End. Bal. $364000

Actual manufacturing overhead= $11000+20000+143000= $174000

Underapplied or overapplied overhead= Actual manufacturing overhead-Applied manufacturing overhead

= $174000-152000= $22000 underapplied

Required 2

Adjusted cost of goods sold= $342000+22000= $364000